A lawyer bills her clients $300 per hour of service. If a client's case requires 17 hours to complete, use a proportion to calculate how much the client will owe the lawyer (excluding tax).
Answer by stanbon(75887)

A lawyer bills her clients $300 per hour of service. If a client's case requires 17 hours to complete, use a proportion to calculate how much the client will owe the lawyer
------------
price/hours = price/hours
----------------------------
x/17 = 300/1
x = 17*300
x = $5100
---------------
P.S. And that is a crime.
Cheers,
Stan H.